OpenStudy (anonymous):

(4)/(x+4)-(7)/(x-2)....i know the answer is (-3)(x+12)/(x+4)(x-2) but what are the steps?

OpenStudy (anonymous):

just take l.c.m and u wil get the answer

OpenStudy (anonymous):

i know the answer i just dont know the steps to get the answer

OpenStudy (anonymous):

apply l.c.m rule

OpenStudy (anonymous):

but then we come up with 28 and its not anywhere near the answer

OpenStudy (dumbcow):

you start w/ different denominators to get a common denominator with binomials multiply them together (x+4)(x-2) For first fraction you added (x-2) so multiply the top by (x-2) 4*(x-2) For 2nd fraction you added (x+4) so multiply top by (x+4) 7*(x+4) this gives 4(x-2) - 7(x+4) all over (x+4)(x-2) distribute on top and add like terms 4x-8-7x-28 =-3x-36 factor out a -3 =-3(x+12) this is still all over (x+4)(x-2) hope that helps

OpenStudy (anonymous):

u wil get it by l.c.m , 4(x-2)-7(x+4) whole divided by (x-2)(x+4) (4x-8-7x-28) whole divided by(x-2)(x+4) (-3x-36) whole divided by (x-2)(x+4) u take -3 common from numerator u wil get u r answer